# A Simple Framework to Accelerate Multilingual Language Model for Monolingual Text Generation

### v1

@@ 4.2 Baselines
We consider the following several baselines for the comparison with the proposed method.
@@
Speculative decoding approach (Chen et al., 2023a; Leviathan et al., 2023) employs a preliminary "draft" model to rapidly generate a set of token candidates at each decoding step. Subsequently, these candidates undergo a validation process by the original language model to ascertain their likelihood as plausible continuations of the text. We implement two variants of this method: one with the capability to reject unsuitable candidates (Spec.) and another without its rejection module (Spec. w/o Rejection). For the draft model, we utilize Llama-2 7B (Touvron et al., 2023). Following the implementation of Chen et al. (2023a), we generate 5 draft tokens at each iteration.
@@
Summarization (0-shot) Translation (3-shot) Method Update Param. Dataset size (Tokens) ROUGE-2 ROUGE-L Morphemes/sec Speed Up BLEU Morphemes/sec Speed Up Vanilla Fine-tuning 13.0B 365M 20.7 36.1 9.8 1.00x 21.2 10.1 1.00x Vocabulary Expansion 13.1B 365M 13.7 23.1 17.1 1.92x 12.3 20.2 2.00x Vocabulary ${}^{\dagger}$ 13.1B 60B + 365M 20.3 37.3 20.5 2.12x 20.3 23.1 2.29x MuMo (Ours) 70M 365M 20.5 36.3 15.3 1.73x 21.7 17.2 1.71x
Table 5: Comparsion with the fine-tuning strategies. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to Vanilla Fine-tuning. Vocabulary ${}^{\dagger}$ was pre-trained on over 60B tokens, comprised of both Korean and English text corpora. Other methods are only trained with the instruction dataset (365M) (Chen et al., 2023c), ShareGPT and Alpaca translated in Korean. The Boldface signifies the superior performances, and the second highest score is underlined.
@@ 5 Further Analysis
Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec BLEU Tokens/sec Ko MuMo 20.3 35.9 55.3 21.7 50.9 w/o Verification 11.0(-9/3) 26.4(-9.5) 60.8(+5.5) 16.3(-5.4) 62.3(+11.4) Ja MuMo 11.6 26.3 59.2 24.3 58.3 w/o Verification 6.7(-4.9) 20.4(-5.9) 69.1(+9.9) 10.8(-13.5) 73.6(+15.3)
@@ 5.1 Comparative Analysis of Fine-Tuning Strategies
In the section, we provide a comparative analysis of three distinct fine-tuning strategies for multilingual models. This analysis aims to highlight the advantages and disadvantages of each strategy, particularly in terms of dataset requirements and the number of parameters to train.
@@ 5.1.1 Setup
1. Vanilla Fine-tuning: This strategy, which serves as a baseline, involves fine-tuning a standard multilingual model on the instruction dataset without any modifications to the pre-trained vocabulary.
@@
To account for the variability of token unit between the different strategies, we report the inference speed with the morphemes per second (Morphemes/sec), providing a standardized measurement.444python-mecab-ko.
@@ 5.2 Initialization of Target Monolingual LM Head
We investigate the impact of 3 different initialization strategies on the target monolingual LM head $g_{\text{mono}}$ in the Target Monolingual LM head. The first strategy involves leveraging embeddings that correspond to the pre-trained representation of a targeted monolingual LM head, termed as Mono-init. The second strategy is initializing the parameters with random value using Gaussian distribution (Random-init). Lastly, we utilize the embeddings from the pre-trained multilingual LM head (Multi-init), as the main experiment. This is achieved by averaging the output embeddings of the multilingual model.
@@
The result demonstrates that the Multi-init approach is almost equally effective with Mono-init. This suggests that our framework can be utilized even without a pre-existing target monolingual model, making it applicable to low-resource languages.
@@ 5.3 Effectiveness of Verification Step
6 Related Work
Tokenization Disparity
Subword tokenization, a common approach in LMs, is typically data-driven. Most of pre-trained tokenizers, which are often trained on predominantly English corpora, frequently result in excessive fragmentation of non-English scripts (Rust et al., 2021; Zhang et al., 2022a). Ahia et al. (2023); Petrov et al. (2023) have found significant tokenization disparities across languages across popular LLMs (Xue et al., 2021, 2022; Scao et al., 2022; OpenAI, 2023) with a depth analysis. Our work endeavors to address the slowdown in inference that arises due to tokenization disparity in non-alphabetic languages.
Modifying Pre-trained Vocabulary
Previous works have explored the adaptation of pre-trained vocabularies or the addition of new tokens (Artetxe et al., 2020; Rust et al., 2021; Liu et al., 2023), these methods often necessitate extensive pre-training to integrate the new tokens effectively (Wang et al., 2019b; Chau et al., 2020; Cui et al., 2023; Liu et al., 2023). In contrast, our MuMo framework sidesteps the need for fine-tuning the parameters of pre-trained models to preserve the original capabilities of the pre-trained language model. Efforts to select items of pre-trained embedding matrix have been made (Abdaoui et al., 2020; Domhan et al., 2022; Ushio et al., 2023), but these have not yielded significant speed up in our experiment, where the relative size of the embedding layer is small (Bogoychev et al., 2023).
Accelerating LLM Inference
The quest to accelerate inference in auto-regressive large language models (LLMs) has led to a variety of approaches. There has been a proliferation of systems specifically engineered for LLM inference, such as Orca (Yu et al., 2022), FlexGen (Sheng et al., 2023), and StreamingLLM (Xiao et al., 2023). Our proposed methodology can be harmonically integrated with the aforementioned techniques. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023a) have also been explored to increase inference velocity. However, the approach often relies on the assumption that a small model can maintain high fidelity when generating a series of multiple tokens. Moreever, acquiring a small yet competitive model may be tricky, especially in a multilingual setup (Conneau et al., 2020; Bandarkar et al., 2023). Our work distinguishes itself by specifically solving the inference inefficiency that arises from excessive fragmentation in the non-alphabetic context.
Parameter Efficient Cross-lingual Transfer Learning
The curse of multilinguality, which refers a trade-off between the language coverage and model capacity (Conneau et al., 2020), is a significant issue even in massively multilingual models, such as mBERT, XLM-R, and mT5 (Devlin et al., 2019; Conneau et al., 2020; Xue et al., 2021; Ansell et al., 2021). The problem has been mitigated through modular parameter-efficient adaptations of the multilingual models through lightweight adapters (Houlsby et al., 2019): additional trainable parameters inserted into the transformer layers of model (Pfeiffer et al., 2020; Üstün et al., 2020; Vidoni et al., 2020; Parović et al., 2022) for a target language. Previous studies focus on improving the representational capability of multilingual models, rather than directly addressing the issue of speeding up multilingual models. In contrast, our research primarily aims to accelerate the inference for multilingual models in specific target languages.
@@ 7 Conclusion
Our study has successfully tackled the challenges in generating text for non-alphabet languages, particularly those associated with excessive fragmentation issues. The approach not only speeds up text generation but also paves the way for more efficient multilingual language applications. Our future work will broaden our experimental scope to languages that were not substantially represented in the pre-trained corpus.
@@ Acknowledgements
We thank Chaehun Park for their valuable discussions and feedback on this paper.
@@
Appendix A Implementation Details
@@
Table 8: Hyperparameters used for generation.

### v2

@@ 4.2 Baselines
We consider the following several baselines for the comparison with the proposed method. Note that all the baselines are implemented instruction-tuned model with multilingual instruction dataset (Chen et al., 2023c).
@@
Speculative decoding approach (Chen et al., 2023a; Leviathan et al., 2023) employs a preliminary "draft" model to rapidly generate a set of token candidates at each decoding step. Subsequently, these candidates undergo a validation process by the original language model to ascertain their likelihood as plausible continuations of the text. We implement two variants of this method: one with the capability to reject unsuitable candidates (Spec.) and another without its rejection module (Spec. w/o Rejection). For the draft model, we utilize Llama-2 7B (Touvron et al., 2023b). Following the implementation of Chen et al. (2023a), we generate 5 draft tokens at each iteration.
@@
  Summarization (0-shot) Translation (3-shot) Method Update Param. Dataset size (Tokens) ROUGE-2 ROUGE-L Morphemes/sec Speed Up BLEU Morphemes/sec Speed Up Vanilla Fine-tuning 13.0B 44M 21.0 36.0 9.8 1.00x 21.4 10.1 1.00x Vocabulary Expansion 13.1B 44M 13.7 23.1 17.1 1.92x 12.3 20.2 2.00x Vocabulary Expansion 13.1B 60B + 44M 20.3 37.3 20.5 2.12x 20.3 23.1 2.29x MuMo (Ours) 70M 44M 20.5 36.3 15.3 1.73x 21.7 17.2 1.71x
Table 4: Comparsion with the fine-tuning strategies. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to Vanilla Fine-tuning. Vocabulary Expansion was pre-trained on over 60B tokens, comprised of both Korean and English text corpora. Other methods are only trained with the instruction dataset (44M tokens) (Chen et al., 2023c), ShareGPT and Alpaca translated in Korean. The Boldface signifies the superior performances, and the second highest score is underlined.
@@ 5 Further Analysis
  Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec BLEU Tokens/sec Ko MuMo 20.3 35.9 55.3 21.7 50.9 w/o Verification 11.0(-9.3) 26.4(-9.5) 60.8(+5.5) 16.3(-5.4) 62.3(+11.4) Ja MuMo 11.6 26.3 59.2 24.3 58.3 w/o Verification 6.7(-4.9) 20.4(-5.9) 69.1(+9.9) 10.8(-13.5) 73.6(+15.3)
@@ 5.1 Comparative Analysis of Fine-Tuning Strategies
In the section, we provide a comparative analysis of three distinct fine-tuning strategies for multilingual models. This analysis aims to highlight the advantages and disadvantages of each strategy, particularly in terms of dataset requirements. and the number of parameters to train.
@@ 5.1.1 Setup
1. Vanilla Fine-tuning: This strategy, which serves as a baseline, involves fine-tuning a standard multilingual model on a target monolingual instruction dataset (44M tokens) without any modifications to the pre-trained vocabulary.
@@
To account for the variability of token unit between the different strategies, we report the inference speed with the morphemes per second (Morphemes/sec), providing a standardized measurement.444python-mecab-ko We only compare the baselines in Korean, because of the availability of model.
@@ 5.2 Initialization of Target Monolingual LM Head
We investigate the impact of three different initialization strategies on the target monolingual LM head $g_{\text{mono}}$ in the Target Monolingual LM head. The first strategy involves leveraging embeddings that correspond to the pre-trained representation of a targeted monolingual LM head, termed as Mono-init. The second strategy is initializing the parameters with random value using Gaussian distribution (Random-init). Lastly, we utilize the embeddings from the pre-trained multilingual LM head (Multi-init), as the main experiment. This is achieved by averaging the output embeddings of the multilingual model.
@@
The result demonstrates that the Multi-init approach is almost equally effective with Mono-init. This suggests that our framework can be utilized some languages that have an off-the-shelf vocabulary set but lack suitable pre-trained representations.
@@ 5.3 Effectiveness of Verification Step
5.4 Comparative Study in Single-Task Training
In the experiment, our primary objective is to investigate whether the inherent capabilities of the instruction-tuned multilingual model, which handles a variety of tasks, could be compromised when trained exclusively on single tasks using either Vocabulary Expansion or MuMo. Both methods introduce newly initialized parameters, raising concerns about potential impacts on the model’s versatility. To address these concerns, we separately trained the model on each task - Question Answering (QA) (Lim et al., 2019; Kurihara et al., 2022) and Summarization (Hasan et al., 2021) - and subsequently conducted a comparative analysis between Vocabulary Expansion and MuMo.
For evaluation, we utilize multiple-task datasets, specifically Korean [^5] Korean-MT-bench and Japanese [^6] Japanese-MT-Bench, which consist solely of questions. For the measurement, We adopt the single-answer grading setup from LLM-as-a-judge (Zheng et al., 2023). This involves presenting a question along with model-generated answers to GPT-4 (acting as the judge) for assessment. The answers are graded on a scale from 1 to 10.
As depicted in Figure 4, the instruction-tuned model initially achieves an average grading of 7.2 in the Korean experiment. However, when fine-tuned using only the QA task, Vocabulary Expansion receives a grading of 1.8, while MuMo receives a grading of 5.9. When trained solely on the summarization task, Vocabulary Expansion receives a grading of 1.6, while MuMo receives a grading of 4.7. Similar trends are observed in the Japanese experiment. The original model receives an average grading of 6.8. When fine-tuned with only the QA task, Vocabulary Expansion receives a grading of 2.1, while MuMo receives a grading of 5.2. When trained exclusively on the summarization task, Vocabulary Expansion receives a grading of 1.2, while MuMo receives a grading of 4.4.
These results suggest that while the grading of the model decreases when trained on single tasks using either method, the decrease is less pronounced with MuMo. This indicates that MuMo is more effective at preserving the model’s multi-task proficiency compared to Vocabulary Expansion. However, it is also clear that neither method can fully maintain the model’s original instruction-following abilities on multiple tasks when trained solely on single tasks. These findings suggest that the instruction dataset, which the model was originally trained on, is crucial for preserving the pre-trained model’s capabilities.
@@ 6 Conclusion
Our study has successfully tackled the challenges in generating text for non-alphabet languages, particularly those associated with excessive fragmentation issues. The approach not only speeds up text generation but also paves the way for more efficient multilingual language applications. Our future work will broaden our experimental scope to languages that were not sufficiently represented in the pre-trained multilingual language model.
Figure 4: Evaluation on multiple-task after training on QA and Summarization task. The red dotted lines represent the average grading of single answers derived from the instruction-tuned multilingual language model. The decline is less pronounced with MuMo, suggesting its relative effectiveness in preserving the model’s multi-task proficiency.
Limitations
Our proposed framework has not been evaluated with languages that exhibit excessive fragmentation issues, such as Tamil, Hebrew, and Arabic (Ahia et al., 2023; Petrov et al., 2023). These languages were not explicitly mentioned in the pre-training corpus of Llama-2 (Touvron et al., 2023b). Additionally, our framework requires off-the-shelf tokenizers for target languages to make Target monolingual LM Head. Our method does not alter the input sequence length, as we focus solely on improving the unit of prediction. This approach This approach differs from the the previous studies (Rust et al., 2021; Cui et al., 2023) which efficiently encode text at the input-level sequence length for excessively tokenized languages. Furthermore, the language models evaluated in the study are restricted to a maximum size of 13B. Larger models, such as Llama-2 30B or 70B, were not implemented due to constraints on available computational resources.
@@ Acknowledgements
We thank Chaehun Park and Daeyoung Kim for their valuable discussions and feedback on the paper.
@@
Appendix A Dataset Details
Training Data
Our study employed a multilingual instruction dataset from Chen et al. (2023c), encompassing Korean and Japanese, for multilingual instruction tuning. Specifically, we utilized ShareGPT and Alpaca-GPT4 for each respective language. The dataset comprises 56k, 55k, and 168k examples for Korean, Japanese and English respectively. To train MuMo LM head, we use ShareGPT and Alpaca-GPT4 (Chen et al., 2023c) in Korean and Japanese for each language.
Evaluation Data
In summarization task, we use validation and test split of XLSum (Hasan et al., 2021), which consist of 1100 examples. We found that more than half of the samples within the validation and test split surpassed the maximum sequence length of Llama-2. Consequently, we filtered out examples exceeding 1536 tokens. From the remaining examples, we randomly selected 300 for our experiments.
Regarding translation task, the dev-test set of FLoRes-200 (Goyal et al., 2022) is employed, consisting of 1012 parallel examples across both languages. We randomly use 3 examples as 3-shot prompts from training set for individual run.
When evaluating multiple-task benchmark dataset 6, we exclude examples in coding and math categories.
Appendix B Additional Results
Experiment on other Language Model
Table 13, and Table 13 present the comparative study in Llama-1 13B (Touvron et al., 2023a) and Mistral 7B Jiang et al. (2023) respectively.
Generation Results
Table 15 and Table 16 present generated texts in summarization and translation tasks.
Appendix C Environment Details
All experiments are implemented using an A100-40GB GPU. The library versions utilized across all experiments include Python 3.9.10, Pytorch 2.1.0, and Transformers 4.34.0.
Appendix D Hyperparameter Details
  Hyperparameter Value Learning rate 2e-5 Epoch 3 Dropout 0.1 Tensor Type bfloat16 Batch size 128 Optimizer AdamW Weight decay 0.01 Warmup ratio 0.04 Maximum sequence length 2048 Learning rate scheduler cosine
Table 7: Hyperparameters settings for multilingual instruction tuning. We follow the script from FastChat Library.
  Hyperparameter Value Learning rate 1e-3 Epoch 3 Dropout 0.1 Tensor Type bfloat16 Batch size 128 optimizer 1.05 Weight decay AdamW Warmup ratio 0.04 Maximum sequence length 2048 Learning rate scheduler cosine $d_{\text{ffn}}$ 1280 non-linearity function $q$ SwiGLU
Table 8: Hyperparameters settings for training MuMo framework.
@@ Appendix D Hyperparameter Details
Table 9: Hyperparameter settings for inference.
Task Evaluation Prompt
Summarization A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. # Document {{sourceDocument}} ## HUMAN: Summarize the document into a {{targetLang}} sentence. ## ASSISTANT:
Translation A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. Translate the following text into {{targetLang}}. ## HUMAN: {{sourceString1}} ## ASSISTANT: {{targetString1}} ## HUMAN: {{sourceString2}} ## ASSISTANT: {{targetString2}} ## HUMAN: {{sourceString3}} ## ASSISTANT: {{targetString3}} ## HUMAN: {{sourceString}} ## ASSISTANT:
Table 10: The evaluation prompt for the main experiment (Sec. 4). We report on 0-shot results on summarization task, and 3-shot results on translation task respectively.
Task Training Prompt
QA A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. # Document {{context}} ## HUMAN: {{question}} ## ASSISTANT: {{answer}}
Summarization A chat between a curious human and an artificial intelligence assistant. The assistant gives helpful, detailed, and polite answers to the human’s questions. # Document {{sourceDocument}} ## HUMAN: Summarize the document into a {{targetLang}} sentence. ## ASSISTANT: {{summary}}
Table 11: The training prompt for the analysis on single-task prompt finetuning (Sec. 5.4).
  Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec Speed Up BLEU Tokens/sec Speed Up Ko Vanilla Decoding 14.7 31.2 29.0 1.00x 18.6 29.7 1.00x MuMo (Ours) 12.8 30.7 45.0 1.51x 18.1 43.0 1.49x Ja Vanilla Decoding 10.4 21.0 28.6 1.00x 20.7 32.6 1.00x MuMo (Ours) 9.6 20.2 54.3 1.89x 20.0 53.8 1.64x
Table 12: Comparative study of the inference speed in Llama-1 13B (Touvron et al., 2023a). The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to the vanilla decoding method.
  Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec Speed Up BLEU Tokens/sec Speed Up Ko Vanilla Decoding 23.0 36.5 34.2 1.00x 18.5 37.3 1.00x MuMo (Ours) 22.8 36.9 56.4 1.65x 18.3 63.2 1.69x Ja Vanilla Decoding 12.9 26.4 33.5 1.00x 27.2 36.8 1.00x MuMo (Ours) 13.2 26.3 64.8 1.93x 26.9 66.3 1.80x
Table 13: Comparative study of the inference speed in Mistral 7B (Jiang et al., 2023). The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to the vanilla decoding method.
Texts (ko) Tokens/sec ROUGE-L
Document 환경부는 22일 사회관계장관회의에서 ’1회용품 함께 줄이기 계획’을 추진한다고 발표했다. 2022년까지 일회용품 사용량을 35% 이상 줄이는 것이 정부의 목표다. 종이 일회용 컵 사용 금지 현재 카페나 빵집 등에서 일회용 플라스틱 컵은 사용이 금지되지만, 종이컵은 사용이 가능했다. 하지만 2021년부터 종이컵 제공 또한 전면 금지된다. 식당, 카페, 급식소에서 플라스틱 빨대, 젓는 막대 등도 2022년부터 금지된다. 매장에서 머그잔에 음료를 받아 마시다 포장해서 가져가려는 경우에도 일회용 컵 사용에 따른 추가 비용을 내야 한다. 환경부는 ’컵 보증금제’ 재도입을 검토 중이다. 소비자가 커피 등 음료를 구매할 때 일정 금액의 보증금을 내고, 컵을 반환하면 그 돈을 돌려받는 방식이다. ’컵 보증금제’는 과거 한 차례 도입됐다가 2008년 폐지됐다. 포장과 음식 배달에서 제공되는 일회용 식기류 무상 제공도 2021년부터 금지된다. 정부는 배달 음식 용기 또한 친환경 소재 또는 다회용기로 전환을 유도하겠다고 발표했다. 장례식장에서도 2021년부터 일회용 식기 용품 사용이 금지된다. 비닐봉지도 금지 현재 비닐봉지는 백화점이나 슈퍼마켓 등 대규모 점포에서는 사용이 금지되어 있다. 편의점 같은 종합 소매업이나 빵집 등에서는 유상으로 구매가 가능하다. 하지만 2022년부터는 제과점이나 가게에서도 일괄 금지된다. 호텔 등 숙박업소의 경우, 50실 이상의 시설에서는 2022년부터 샴푸, 린스, 칫솔 등 일회용 위생용품 무상 제공이 금지된다. 2024년부터는 모든 숙박업소에 일괄 적용된다. 택배 포장재 줄이기 최근 택배와 신선식품 배송이 급격히 늘면서, 환경부는 배송용 포장재 사용량 증가 해소를 위한 사업도 추진한다고 발표했다. 과대포장 문제가 제기된 배송 상품의 경우 포장기준을 강화하고, 업계와 협의해 종이 완충재와 테이프 없는 상자 등 친환경 포장재를 마련할 계획이다. 2020년부터 이미 포장된 제품을 다시 포장해서 묶어 판매하는 소위 이중 포장 행위가 금지된다. 올해 13살인 라니엘은 8살 때부터 강물에 떠내려온 쓰레기를 줍기 시작했다 다른 나라는 2018년 10월, 유럽연합은 바다 오염을 막기 위해 일련의 일회용 플라스틱 제품 사용을 완전히 금하는 법안을 통과시켰다. 유럽연합은 2021년부터 법안이 발효할 것으로 기대하고 있다. 금지 품목에는 플라스틱 식기류, 빨대, 면봉 등이 있으며 식품과 음료에 사용되는 플라스틱 컵 등 일회용 플라스틱 사용 역시 줄이도록 하는 내용을 담았다. 인도의 경우 2022년부터 일회용 플라스틱 사용이 전면 금지된다.
GT 2021년부터 카페에서 음료를 포장할 경우, 일회용 컵을 무상으로 사용하지 못한다.
Vanila Decoding 환경부는 2022년까지 일회용품 사용량을 35% 이상 줄이는 것을 목표로 ’일회용품 함께 줄이기 계획’을 추진한다고 발표했다. 27.7 33.4
MuMo 환경부에 따르면 2022년까지 일회용품 사용량에서 35% 이상 줄이기를 목표로 하며, 현재는 일회용 플라스틱 컵 사용이 금지되었으며 2021년부터는 종이컵 제공도 금지될 예정입니다. 47.2 38.1
Table 14: Generated texts on summarization task in Korean. The sample is extracted from the validation set of XLSum (Hasan et al., 2021). GT indicates the ground truth summary of the example.
Texts (ja) Tokens/sec ROUGE-L
Document {CJK} UTF8min 犬のマックスは16時間、女の子に寄り添った 女の子のオーロラちゃんは前の日から行方が分からなくなり、家族や警察など約100人が捜索に当たっていた。 クイーンズランド・サザンダウンズの自宅を出て原野に迷い込んだオーロラちゃんの後を、犬のマックスが追い、16時間近くずっと寄り添っていたとみられている。高齢のマックスは、目と耳が部分的に不自由。 1人と1匹が丘の斜面で一緒にいるのを、親族が21日朝に発見した。 オーロラちゃんの祖母、レイサ・マリー・ベネットさんは、自宅から約2キロ離れた場所で、オーロラちゃんの叫び声を聞いたと豪ABCに話した。 「大急ぎで山を駆け上がって上までたどりつくと、犬がこちらに向かってきて、オーロラのところへ一直線に連れて行ってくれた」 親類によると、気温が15度まで下がるなか、オーロラちゃんは犬のマックスと岩の下に避難していたという。 警察車両の横に立つマックス。動物の専門家によると、高齢な犬ほど人間とのつながりを特に重視するという クイーンズランド警察は、マックスの行動を称え、名誉警察犬の地位を与えた。 「3歳の子供なら、夜間とても怖かっただろうし、とても寒かったはずだ。犬が寄り添っていたおかげで、女の子は心強かっただろうし、寒くならなかったのだろう。明るい結末でよかった」とクレイグ・ベリー警部は話した。 ツイッターでは大勢が、マックスをほめちぎり、おやつをたくさんあげてほしいと書き込んだ。 なぜずっとそばに なぜマックスがずっとオーロラちゃんのそばを離れなかったのかについて、シドニー大学のポール・クリービー教授(動物行動学)は、高齢な犬ほど人間とのつながりを大事にするので、女の子の動揺を察知したのだろうと話す。 「もし女の子が泣いていたなら、犬は元気付ける行動をとった可能性が高い」とクリービー教授はBBCに話した。 「女の子のそばにずっといて、支えてあげるのが、なにより大事な行動だったはずだ」
GT {CJK} UTF8min 豪クイーンズランドの警察は21日午前、原野に迷い出て行方不明になった3歳少女を発見したと発表すると共に、家族の17歳になる牧牛犬が女の子に約16時間寄り添っていたと明らかにした。
Vanila Decoding {CJK} UTF8min クイーンズランド州のサザンダウンズに住む小さな女の子オーロラが、家族や警察など約100人が捜索に当たっていたが、16時間前に行方不明になった。 29.2 27.8
MuMo {CJK} UTF8min オーストラリアのクイーンズランド州サバーンダウンズ地域で行方不明になった子供を見つけ、オーストラリア警察の名誉警察犬になった高齢の犬マックス(Max)についてのジャーナル記事とともに、オーストラリア警察の名誉警察犬になったという内容です。 57.3 31.9
Table 15: Generated texts on summarization task in Japanese. The sample is extracted from the validation set of XLSum (Hasan et al., 2021). GT indicates the ground truth summary of the example.
[…]
Source Since moving to the Catalan-capital, Vidal had played 49 games for the club.
GT 바르셀로나로 이적한 후 비달은 클럽을 위해 49경기를 뛰었습니다.
Vanila Decoding 바르셀로나로 이적한 이후로 비달은 이 클럽에서 49경기에 출전했습니다. 27.2 22.8
MuMo 바르셀로나로 이적했던 비달은 클럽에서 총 49경기를 출전했습니다. 45.7 26.9
[…]
Source Just after 11:00, protesters blocked traffic on the northbound carriage in Whitehall.
GT 11시가 막 지난 후, 시위대는 화이트홀에 있는 북쪽으로 향하는 마차들의 교통을 막았다.
Vanila Decoding 백알 화이트홀에서 오전 11시 15분경, 시위자들이 북쪽 차선을 차단하여 교통을 방해했습니다. 27.6 5.1
MuMo 백알 화이트홀에서 오후 11시 이후, 시위대가 북쪽 선행 차량을 차단했습니다. 44.3 4.3
[…]
Source Since moving to the Catalan-capital, Vidal had played 49 games for the club.
GT {CJK} UTF8minカタルーニャの州都に移って以来、ビダルはクラブで49試合に出場しました。
Vanila Decoding {CJK} UTF8min バルセロナに移動してから、ビダルさんは約49試合でプレーしています。 27.5 6.8
MuMo {CJK} UTF8min バルセロナに移動してから、ビダルさんは約49試合でプレーしています。 52.2 6.8
[…]
Source Just after 11:00, protesters blocked traffic on the northbound carriage in Whitehall.
GT {CJK} UTF8min 11時すぎちょうどに抗議者たちはホワイトホールの北行き車両の交通を遮断しました。
Vanila Decoding {CJK} UTF8min 11時過ぎに、ホワイトホールの北行線上で抗議者が交通を妨害しました。 28.3 26.7
MuMo {CJK} UTF8min 午前11時過ぎ、デモ隊はホワイトホールの北へ向かう馬車の交通を阻止した。 50.3 25.0
Table 16: Generated texts on translation task. The samples are extracted from the dev-test set of FLoRes-200 (Goyal et al., 2022). GT indicates the ground truth sentence of the example.
  Summarization (0-shot) Translation (3-shot) LM Head Initialization ROUGE-2 ROUGE-L BLEU Mono-init 20.7 36.2 21.5 Random-init 19.2 35.5 17.2 Multi-Init 20.3 36.3 21.7 Focus-Init 20.8 36.5 21.9
Table 17: Comparative analysis for the initialization strategy We exploit FOCUS (Dobler and de Melo, 2023) embedding to initialize the Target monolingual LM Head. Our framework can be harmonically integrated with the initialization strategy of multilingual token embedding.